1. For u = (3,-2,-2) and v = (1, 2, 1), find a unit vector parallel to 3u +2v.

Please show all work and steps I will provide like for answer
Answer #1

34+ 2V = 3 (3,-2, -22 +2(1,2,1) = (9,-6,-6)+(2,4,2) (11, -2,-4) parallel to unit veetor will be V121+4+16 ( 11, -2,-4) (0:27